# Knowledge-to-SQL: Enhancing SQL Generation with Data Expert LLM

### v2

@@ Abstract
Generating accurate SQL for user queries (text-to-SQL) is a long-standing problem since the generation of the SQL requires comprehending the query and database and retrieving the accurate data from the database accordingly. Existing models rely on the comprehensive ability of Large Language Models (LLMs) to generate the SQL according to the database schema. However, there is some necessary knowledge that is not explicitly included in the database schema or has been learned by LLMs. Thus, the generated SQL of the knowledge-insufficient queries may be inaccurate, which negatively impacts the robustness of the text-to-SQL models. To deal with this situation, we propose the Knowledge-to-SQL framework, which employs tailored Data Expert LLM (DELLM) to provide helpful knowledge for all types of text-to-SQL models. Specifically, we provide the detailed design of DELLM, in terms of table reading, and the basic fine-tuning process. We further provide a Preference Learning via Database Feedback (PLDBF) training strategy to guide the DELLM to generate more helpful knowledge for LLMs. Extensive experiments verify DELLM can enhance the state-of-the-art LLMs on text-to-SQL tasks. The model structure and the parameter weight of DELLM are released for further research.
@@ 1 Introduction
Generating SQL based on user queries (Text-to-SQL) is currently one of the leading real-world applications for Large Language Models (LLMs). The most straightforward approach is to input the user queries and database schema into the LLMs and rely on them to generate the SQL Dou et al. (2022); Liu et al. (2023a); Pourreza and Rafiei (2023). However, in real-world applications, user queries and database organization may contain customized or specialized knowledge Yu et al. (2018); Li et al. (2023b). In such cases, the intuitive approach may result in inaccurate or un-executable SQL unless a human expert provides the necessary knowledge to the LLMs Li et al. (2023b). Given this challenge, it is valuable to develop a non-human architecture that can automatically generate the required data expert knowledge for SQL generation. This would significantly enhance the performance and robustness of text-to-SQL models.
Figure 1: A sketch map illustrating the significance of incorporating expert knowledge in Text-to-SQL task.
@@
Existing models primarily focus on fully exploring the comprehensibility of Language Model. As the forerunner, T5-based methods Scholak et al. (2021); Li et al. (2023a); Rai et al. (2023) made an initial attempt to train a language model from scratch to generate SQL based on user queries and database schemas. Following the popularity of Language Models, particularly ChatGPT and GPT4 OpenAI (2023), DIN-SQL Pourreza and Rafiei (2023) utilizes GPT to decompose the process of SQL generation into several sub-tasks. It uses GPT to first accomplish these sub-tasks and then aggregates the results to generate the SQL. Similarly, DAIL-SQL Gao et al. (2023) refers to SQLs from similar queries to accomplish the generation of the required SQL. Most recently, MAC-SQL Wang et al. (2023) assigns different roles to three GPT4 agents, specifically Selector, Inspector, and Refiner, and uses these agents to contribute to the SQL generation process.
@@
However, existing models mainly focus on designing more complex SQL generation structures in order to improve SQL generation performance. However, there is a lack of emphasis on the necessity of data expert knowledge Li et al. (2023b). As depicted on the left of Figure 1, relying solely on the user’s query and the database schema without the data expert knowledge may generate inaccurate conditions. For example, the incorrect condition “y.Date like ”2012-08-%”” could potentially result in the SQL returning empty data. In contrast, on the right of Figure 1, when provided with accurate data expert knowledge, the LLMs are able to rectify the condition and generate the valid SQL.
@@
Nevertheless, generating expert knowledge from data faces the following challenges: 1. Query & Database Specialization: The generated expert knowledge should be specialized for the given query and database. It is challenging for the knowledge generator to understand the query and database and provide suitable knowledge. 2. Table Content Awareness Cheng et al. (2023): The expert should be able to read the content of the table in order to determine whether it is necessary to provide detailed content examples in the knowledge. 3. Performance Enhancement Li et al. (2023b): The generated knowledge should be helpful for the SQL generation models. It is challenging to ensure that the expert knowledge can contribute to more accurate SQL generation.
@@
By addressing the aforementioned challenge, we present a complete Knowledge-to-SQL pipeline for enhancing the SQL generation of LLMs. Specifically, we propose a detailed design for the Data Expert Language Model (DELLM), which comprises of a table reading component and a knowledge-oriented fine-tuning mechanism. Furthermore, we introduce Preference Learning via Database Feedback (PLDBF) to further improve the helpfulness of the generated data expert knowledge in LLM-based SQL generation Christiano et al. (2017). Specifically, PLDBF provides rewards to DELLM based on two tailored criteria: 1) the extent to which the generated knowledge aids in retrieving more accurate data, and 2) the degree to which the generated knowledge assists LLM in producing more precise SQL queries. In summary, our contributions can be listed as follows:
@@
  • We highlight the significance of data expert knowledge and present a comprehensive Knowledge-to-SQL pipeline for improving SQL generation.
@@
  • We introduce a specially designed Data Expert Language Model (DELLM), along with customized structure, fine-tuning technique, and preference-tuning training strategies.
@@
  • We release the training and evaluation code of DELLM as open source, and provide parameter weights for future research.
@@
  • We validate the effectiveness of our approach on the Bird and Spider datasets, demonstrating that DELLM can generally enhance the performance of common text-to-SQL models.
@@ 2.1 LLMs for Text-to-SQL
The Text-to-SQL task focuses on translating natural language queries into SQL statements. Recent developments in this field have shown a growing interest in using large language models (LLMs) paired with prompts. Chain of Thoughts (CoT) Wei et al. (2022b), an effective prompt engineering technique, has found considerable utility in the Text-to-SQL domain. Numerous researchers conducted empirical studies on example selection and prompt organization for Text-to-SQL using CoT Rajkumar et al. (2022); Gao et al. (2023); Chang and Fosler-Lussier (2023); Li et al. (2023b). In DAIL-SQL Gao et al. (2023), the authors introduced a novel approach for selecting pertinent examples, considering both the user’s query and SQL statement similarities. Moreover, ACT-SQL  Zhang et al. (2023a) employed a cost-efficient method to automatically generate CoT instances, alleviating the need for manual prompt creation. For enhancing the reasoning capabilities of LLMs, works like C3  Dong et al. (2023c) and DIN-SQL Pourreza and Rafiei (2023) have proposed decomposing main tasks into multiple sub-tasks, which have been specifically designed for zero-shot and few-shot Text-to-SQL tasks respectively. Most recently, MAC-SQL Wang et al. (2023) assigns multiple GPT4 agents with different roles for the SQL generation process. Additionally, to assist Text-to-SQL with external knowledge, authors in Li et al. (2023b) engaged human experts to annotate relevant knowledge for each Text-to-SQL instance. However, this mode of human-led annotation proves highly expensive for large-scale tasks. In this paper, we focus on exploring automated knowledge generation for text-to-SQL using LLMs.
@@ 2.2 Prompt Engineering in Text-to-SQL
Previous studies have highlighted the importance of prompt engineering in optimizing the performance of LLMs Radford et al. (2019); Liu et al. (2023b). The Effective prompt design significantly improves the efficiency and quality of LLM outputs Wei et al. (2022a). Research initiatives like Chain of Thoughts Wei et al. (2022b) and Retrieval Augmented Generation Lewis et al. (2020) integrate contextual information to enhance LLMs’ understanding of natural language reasoning and comprehension. In the text-to-SQL task, recent investigations Rajkumar et al. (2022) have refined the prompts for LLMs, resulting in high uniformity input formats. These prompts typically include user queries, database schemas, and text-to-SQL instructions. However, comprehending lengthy input schemas and correlating them with user queries poses a significant challenge for LLMs, stemming from diverse aspects of prompt design. Leading approaches Rajkumar et al. (2022); Pourreza and Rafiei (2023); Zhang et al. (2023a) have incorporated techniques such as few-shot examples Gao et al. (2023) and chain of thoughts Li et al. (2023b); Wang et al. (2023) guidelines to address this challenge. Our task can also be viewed as a particular prompt engineering, with the purpose of generating data expert knowledge as input prompt to assist LLMs understanding of text-to-SQL.
3 Purposed Method
Figure 2: The overview of our approach. Upper is the overall knowledge-to-SQL framework. The details of DELLM are presented at the bottom. On the left side, we have the framework of DELLM, and on the right side, we introduce Preference Learning via Database Feedback (PLDBF), which is employed to enhance the performance of DELLM.
We adhere to the two criteria mentioned to guide knowledge generation: 1) Enhancing precise database execution. 2) Improving precise SQL generation. Our proposed approach aims to develop a model that satisfies the above criteria. As depicted in Figure 2, our methodology consists of the following three primary modules:
  • The supervised fine-tuned (SFT) model: It takes the user query, schema, and relevant table readings from the database as input and converts them into generated knowledge.
  • The preference learning (PL) framework: It enhances the model by aligning the feedback from executing queries on the database with the contribution of ground-truth SQL.
  • An off-the-shelf text-to-SQL model: It predicts the SQL by inputting the query, schema, and generated knowledge.
@@ 3.1 Supervised Fine-Tuning of DELLM
This module aims to generate knowledge based on the user query and the database schema. Assuming the user query is $Q$ and the corresponding database schema is $\mathcal{S}$, our goal is to match the relevant sub-table $\mathcal{T}_{\alpha}$ and generate knowledge $K^{gen}$ with the aforementioned inputs.
Table Reading
In our study, we incorporate the task of table reading to generate relevant knowledge. However, when dealing with large databases, inputting the complete database tables poses a challenge in terms of input length and the inclusion of redundant information. To address this, we utilize semantic techniques to match the target table. This allows us to extract the most pertinent table for the given question as the input prompt for the Language Models (LLMs). Let $\mathcal{S}$ represent the database schema, it can be defined as:
@@
Where $\mathcal{T}_{i}$ denotes the $i^{th}$ table in the database, and $c_{ij}$ represents the $j^{th}$ column in table $\mathcal{T}_{i}$. Given a schema with $n$ tables, we can denote the collection of each table’s column as $\mathcal{C}=\{\{c_{11},...,c_{1m_{1}}\},...,\{c_{n1},...,c_{nm_{n}}\}\}$. For each column $c_{ij}\in\mathcal{C}$, we can obtain a sub-collection of relevant columns denoted as $\mathcal{C}_{\alpha}$:
@@
where $\operatorname{sim}(\cdot)$ denotes the semantic similarity calculator, and $\operatorname{sim}(Q,c_{ij})>\alpha$ represents the similarity between column $c_{ij}$ and query $Q$ exceeding the threshold $\alpha$. The reading tables can be given as:
@@
Supervised Fine-tuning
We rephrase the input prompt of our knowledge generation model, and then the generation process can be represented as:
@@
where $\pi(\cdot)$ represents a text generation model, which we utilize to generate the knowledge $K^{gen}$ as defined earlier. During this training phase, the objective loss function can be defined as:
@@
where $K^{gold}$ denotes the ground-truth knowledge annotated by human experts. The ultimate goal is to minimize the aforementioned cross-entropy loss by means of supervised fine-tuning. Consequently, we can acquire a supervised fine-tuned model $\pi^{SFT}$, which possesses the ability to generate knowledge in a preliminary manner.
@@ 3.2 Feedback From Database Execution and Ground-truth SQL Contribution
To make our model capable of generating knowledge to assist accurate database execution and SQL contribution, we use preference learning to enhance the ability through feedback from the database’s execution and ground-truth SQL contribution.
Feedback From Database Execution
In section 3.1, we obtained a model $\pi^{\text{SFT}}$ via supervised fine-tuning. We utilize the knowledge generated by this model to aid an off-the-shelf text-to-SQL model $g(\cdot)$ in SQL generation.
@@
where $\hat{Y}^{gen}$ represents the SQL statement generated using the knowledge $K^{gen}$. For every instance in the training set, we obtain a predicted SQL set $\mathcal{\hat{Y}}^{gen}$. Similarly, we generate a predicted SQL set $\mathcal{\hat{Y}}^{gold}$ using the manually annotated ground-truth knowledge $K^{gold}$. To interact with the database execution, we execute the SQL sets $\mathcal{\hat{Y}}^{gen}$ and $\mathcal{\hat{Y}}^{gold}$, respectively, to obtain the result sets $\hat{V}^{gen}$ and $\hat{V}^{gold}$. An indicator function is defined to evaluate the execution results. We have:
@@
then, we utilize this indicator function as the condition for annotating the preference knowledge pair in the feedback of database execution.
$\mathcal{P}^{db}_{\{K_{w},K_{l}\}}=\{K^{gold}_{i},K^{gen}_{i}\mid\mathds{1}_{db}(\hat{V}^{gold}_{i},\hat{V}^{gen}_{i})=0\}.$ (8)
@@
Feedback From Ground-truth SQL Contribution
@@
Table 1: Example of ground-truth SQL contribution. We highlight the key content in SQL and the corresponding contributing and non-contributing knowledge.
Knowledge plays a decisive role in text-to-SQL, that the final predicted SQL will include sub-knowledge information introduced by the given knowledge. An example is shown in Table 1. However, the incorrect sub-knowledge may lead to a misprediction in SQL statements. A contributing knowledge indicates that every sub-knowledge can correctly guide the SQL prediction. Thus, we focus on studying the contribution of generated knowledge to the ground-truth SQL.
@@
where $Y$ is the ground-truth SQL. We check whether every sub-knowledge contains $Y$ to judge whether the knowledge contributes. Then, we can obtain the preference dataset of ground-truth SQL contribution feedback by:
@@
This feedback necessitates the generated knowledge to become SQL contributing and reduce the redundancy in the knowledge generation process.
By Eq. 8 and 10, we obtain two preference set $\mathcal{P}^{db}$ and $\mathcal{P}^{sql}$. To get the final preference learning dataset $\mathcal{D}$, we combine corresponding input of $K_{l}$ according to Eq. 4, denoted by $I_{l}$ and the preference pair $(K_{w},K_{l})$ by:

### v3

@@ Abstract
Generating accurate SQL queries for user questions (text-to-SQL) has been a long-standing challenge since it requires a deep understanding of both the user’s question and the corresponding database schema in order to retrieve the desired content accurately. Existing methods rely on the comprehensive capability of large language models (LLMs) to generate the SQL. However, some necessary knowledge is not explicitly included in the database schema and user question or has been learned by LLMs. Thus, the generated SQL of the knowledge-insufficient questions may be inaccurate, negatively influencing the text-to-SQL models’ performance and robustness. To address this challenge, we propose the Knowledge-to-SQL framework, which employs tailored Data Expert LLM (DELLM) to provide helpful knowledge for all text-to-SQL models. Specifically, we introduce the detailed implementation of DELLM regarding table reading and the basic fine-tuning process. We further propose a Preference Learning via Database Feedback (PLDBF) strategy, refining the DELLM to generate more helpful knowledge for LLMs. Extensive experiments verify that DELLM can enhance the state-of-the-art approaches for text-to-SQL tasks. The corresponding code of DELLM is released for further research [^1] https://github.com/Rcrossmeister/Knowledge-to-SQL.
@@ 1 Introduction
Generating SQL based on user questions (text-to-SQL) is currently one of the leading applications for large language models (LLMs). The most straightforward approach is to input the user questions and database schema into the LLMs and rely on their capability of natural language understanding to generate the SQL Dou et al. (2022); Liu et al. (2023a); Pourreza and Rafiei (2023). However, in real-world applications, user queries and database structure may contain customized or specialized knowledge Yu et al. (2018); Li et al. (2023b), including arithmetic reasoning, domain knowledge, synonym explanation, etc. Since the knowledge is not explicit in either the user question or the database schema Dou et al. (2022); Gan et al. (2021). In such cases, the intuitive approach may result in inaccurate or un-executable SQL unless a human expert provides the necessary helpful knowledge (so-called “expert knowledge”) to the LLMs to bridge the knowledge gap with database content Li et al. (2023b); Gan et al. (2021). Given this challenge, it is valuable to develop a non-human data expert system that can automatically generate the required expert knowledge to assist SQL generation. This would significantly enhance the performance and robustness of text-to-SQL implementations.
Figure 1: A sketch map illustrating the significance of incorporating expert knowledge in the text-to-SQL implementation. In the given example, the generation without expert knowledge makes mistakes in arithmetic reasoning and data conditions. Expert knowledge bridges the knowledge gap between the LLMs and the database, which assists the LLMs in generating accurate SQL.
@@
Existing methods primarily focus on fully exploring the comprehensibility of the pre-trained language models (PLMs). As the forerunner, T5-based methods Scholak et al. (2021); Li et al. (2023a); Rai et al. (2023) initially attempted to train PLMs from scratch to generate SQL based on user questions and database schema. Following the popularity of proprietary LLMs, particularly ChatGPT and GPT-4 OpenAI (2023), DIN-SQL Pourreza and Rafiei (2023) utilizes LLMs to decompose the process of SQL generation into several sub-tasks. It uses LLMs to first accomplish these sub-tasks and then aggregates the results to generate the final SQL. Similarly, DAIL-SQL Gao et al. (2024) designed a strategy for sampling few-shot instances, then using few-shot prompting for proprietary LLMs to accomplish the generation of the required SQL. Most recently, MAC-SQL Wang et al. (2023) assigns different roles to three LLM agents, specifically selector, decomposer, and refiner, then uses these agents to make different contributions to the text-to-SQL process.
@@
However, recent methods mainly focus on designing more sophisticated structures in order to improve text-to-SQL performance. Although achieving promising performance, there is a lack of emphasis on the necessity of expert knowledge Li et al. (2023b). As depicted on the left of Figure 1, relying solely on the user’s question and the database schema as input without expert knowledge may cause inaccurate conditions in the generated SQL. For example, the incorrect condition “y.Date like ”2012-08-%”” could potentially result in the SQL returning empty data. In contrast, on the right of the figure, when provided with accurate expert knowledge for assistance, the LLMs are able to rectify the condition and generate valid SQL.
@@
Nevertheless, generating expert knowledge from the question and schema faces the following challenges: 1. Question & Database Specialization: The generated expert knowledge should be specialized for the given question and database. It is challenging for the data expert to understand the question and database and provide helpful knowledge. 2. Table Content Awareness Cheng et al. (2023): The data expert should be able to read the content of the table in order to determine whether it is necessary to provide detailed content examples in the knowledge. 3. Performance Enhancement Li et al. (2023b): The generated knowledge should be helpful for the text-to-SQL models. It is challenging to ensure that expert knowledge can contribute to more accurate SQL generation.
@@
By addressing the aforementioned challenges, we present a detailed design of Knowledge-to-SQL framework for enhancing the SQL generation of LLMs. Specifically, we propose a well-designed Data Expert Large Language Model (DELLM), which comprises a table reading module and a knowledge-oriented supervised fine-tuning process. Furthermore, we introduce Preference Learning via Database Feedback (PLDBF) to further refine the helpfulness of the generated expert knowledge for LLM-based text-to-SQL Christiano et al. (2017); Rafailov et al. (2023); Hong and Liu (2024). Specifically, PLDBF provides preferences to DELLM based on two tailored criteria: 1) the extent to which the generated knowledge aids in retrieving more accurate content, and 2) the degree to which the generated knowledge assists LLM in producing more precise SQL queries. In summary, our contributions can be listed as follows:
@@
  • We highlight the significance of expert knowledge and present the knowledge-to-SQL framework for improving SQL generation.
@@
  • We introduce a well-designed Data Expert Large Language Model (DELLM), along with customized structure, fine-tuning technique, and preference-tuning training strategies.
@@
  • We release the training and evaluation code of DELLM as open source for future research.
@@
  • We validate the effectiveness of our approach on the BIRD and Spider datasets, demonstrating that DELLM can generally enhance the performance of common LLM-based text-to-SQL implementations.
@@ 2.1 LLMs for Text-to-SQL
The text-to-SQL task focuses on translating natural language questions into SQL queries. Recent advances in this field have shown a growing interest in using large language models (LLMs) paired with prompt engineering. Chain of Thought (CoT) prompting Wei et al. (2022b), an effective prompt engineering technique, has found considerable utility in the text-to-SQL domain. Numerous researchers conducted empirical studies on prompt organization for text-to-SQL using CoT Rajkumar et al. (2022); Gao et al. (2024); Chang and Fosler-Lussier (2023); Li et al. (2023b); Zhang et al. (2023a). In DAIL-SQL Gao et al. (2024), the authors introduced a novel approach for selecting pertinent few-shot examples, considering both similarities in the user’s question and SQL query. Moreover, ACT-SQL Zhang et al. (2023a) employed a cost-efficient method to automatically generate CoT instances, alleviating the need for manual prompt creation. For enhancing the reasoning capabilities of LLMs, works like C3  Dong et al. (2023b) and DIN-SQL Pourreza and Rafiei (2023) have proposed a paradigm for decomposing main tasks into multiple sub-tasks, which have been specifically designed for zero-shot and few-shot text-to-SQL tasks respectively. Most recently, MAC-SQL Wang et al. (2023) assigns multiple LLM agents with different roles for the text-to-SQL process. Additionally, to assist text-to-SQL with expert knowledge, authors in Li et al. (2023b) engaged human experts to annotate helpful knowledge for each text-to-SQL instance. However, this mode of human-led annotation proves to be highly labor-consuming for large-scale tasks. In this paper, we focus on exploring automated expert knowledge generation for assisting text-to-SQL using LLMs.
@@ 2.2 Prompt Engineering in Text-to-SQL
Previous studies have highlighted the importance of prompt engineering in optimizing the performance of LLMs Radford et al. (2019); Liu et al. (2023b). The effective prompt design significantly improves the efficiency and quality of LLM outputs Wei et al. (2022a). Research initiatives like CoT Wei et al. (2022b) and retrieval augmented generation (RAG) Lewis et al. (2020) integrate contextual information to enhance LLMs’ understanding of natural language reasoning. In the text-to-SQL task, recent investigations Rajkumar et al. (2022) have refined the prompts for LLMs, resulting in high uniformity input formats. These prompts typically include user questions, database schema, and task instructions. However, comprehending complex input schema and correlating them with user questions poses a significant challenge for LLMs, stemming from diverse aspects of prompt design. Leading approaches Rajkumar et al. (2022); Pourreza and Rafiei (2023) have incorporated techniques such as few-shot sampling Gao et al. (2024) and CoT Li et al. (2023b); Wang et al. (2023) guidelines to address this challenge. Our proposed framework can also be viewed as a particular prompt engineering, with the purpose of generating expert knowledge as input prompts to assist LLMs’ understanding of text-to-SQL.
3 Proposed Method
Figure 2: The overview of our approach. The upper is the overall knowledge-to-SQL framework. The details of DELLM are presented at the bottom. On the left side, we have the framework of DELLM, including supervised fine-tuning (SFT) and table reading. On the right side, we introduce preference learning via database feedback (PLDBF), which is employed to further refine the performance of DELLM.
We adhere to the two criteria mentioned to guide knowledge generation: 1) Enhancing accurate database execution. 2) Improving precise SQL generation. Our proposed method aims to develop a system called DELLM for expert knowledge generation that satisfies the above criteria. As depicted in Figure 2, our framework consists of the following three primary modules:
  • The supervised fine-tuned (SFT) model: It takes the user question, database schema, and relevant tables from the database as input and converts them into generated knowledge.
  • The preference learning (PL) framework: It refines the model by aligning the feedback from SQL query executions on the database with the contributions from ground-truth SQL.
  • An off-the-shelf text-to-SQL model: It predicts the SQL by inputting the user question, database schema, and generated knowledge.
@@ 3.1 Supervised Fine-Tuning of DELLM
This module aims to generate expert knowledge based on the user question and the database schema. Assuming the user question and the corresponding database schema are $Q$ and $\mathcal{S}$ respectively; the goal is to match the relevant sub-tables $\mathcal{T}_{\alpha}$ and generate knowledge $K^{gen}$ with the above inputs.
Table Reading.
In our study, we incorporate the task of table reading to generate expert knowledge. When dealing with large databases, inputting the complete database tables poses a challenge regarding input length limitation and redundancy. To address this, we utilize semantic techniques to match the relevant table. This allows us to extract the most pertinent table for the given question as the input prompt for the subsequent SFT model. Let $\mathcal{S}$ represent the database schema, defined as:
@@
where $\mathcal{T}_{i}$ denotes the $i^{th}$ table in the database, and $c_{ij}$ represents the $j^{th}$ column in table $\mathcal{T}_{i}$, $m_{i}$ denotes the number of column included in $\mathcal{T}_{i}$. Given a schema with $n$ tables, we can denote the collection of all table’s columns as $\mathcal{C}=\{\{c_{11},...,c_{1m_{1}}\},...,\{c_{n1},...,c_{nm_{n}}\}\}$. For each column $c_{ij}\in\mathcal{C}$, we can obtain a sub-collection of relevant columns denoted as $\mathcal{C}_{\alpha}$:
@@
where $\operatorname{sim}(\cdot)$ denotes the semantic similarity calculator, and $\operatorname{sim}(Q,c_{ij})>\alpha$ represents the similarity between column $c_{ij}$ and query $Q$ exceeding the threshold $\alpha$. The relevant sub-tables can be obtained by:
@@
Supervised Fine-tuning.
The knowledge generation process can be represented as follows:
@@
$\pi(\cdot)$ represents a text generation model we utilize as a backbone model for the SFT process. As defined earlier, the model is required to generate the knowledge $K^{gen}$. The objective function of SFT can be defined as:
@@
where $K^{gold}$ denotes the gold (ground-truth) knowledge annotated by human experts. The SFT process aims to minimize the cross-entropy loss in Eq. 5 defined above. Consequently, we can obtain a fine-tuned model $\pi^{SFT}$, which possesses the ability to generate knowledge in a preliminary manner.
@@ 3.2 Feedback From Database Execution and Ground-truth SQL Contribution
To enable our model to generate helpful knowledge for accurate database execution and contribute effectively to SQL queries, we employ the preference learning framework. This framework refines the model’s capabilities by aligning feedback from database executions and the contributions of ground-truth SQL.
Feedback From Database Execution.
In Section 3.1, we obtained a model $\pi^{\text{SFT}}$ via SFT. We utilize the knowledge generated by this model to assist an off-the-shelf text-to-SQL model $g(\cdot)$ in SQL generation.
@@
where $\hat{Y}^{gen}$ represents the SQL query generated using the knowledge $K^{gen}$. For every instance in the training set, we obtain a predicted SQL; the collection of these SQL queries is denoted by $\mathcal{\hat{Y}}^{gen}$. Similarly, we obtain a predicted SQL collection $\mathcal{\hat{Y}}^{gold}$ using the manually annotated ground-truth knowledge $K^{gold}$. To interact with the database execution, we execute the SQL collections $\mathcal{\hat{Y}}^{gen}$ and $\mathcal{\hat{Y}}^{gold}$, respectively, to obtain the result sets $\hat{V}^{gen}$ and $\hat{V}^{gold}$. An indicator function is defined to evaluate the execution results. We have:
@@
then, we utilize this indicator function as the condition for annotating the preference knowledge pairs with the feedback of database execution:
$\mathcal{P}^{db}_{\{K_{w},K_{l}\}}=$ (8)
$\{K^{gold}_{i},K^{gen}_{i}\mid\mathds{1}_{db}(\hat{V}^{gold}_{i},\hat{V}^{gen}_{i})=0\}.$
@@
Feedback From Ground-truth SQL Contribution.
@@
Table 1: Example of ground-truth SQL contribution. We highlight the key content in SQL and the corresponding sub-knowledge included in the contributing and non-contributing knowledge.
Knowledge plays a decisive role in text-to-SQL that the final predicted SQL will include sub-knowledge content introduced by the given knowledge. An example is shown in Table 1. However, the incorrect sub-knowledge may lead to a misprediction in generating SQL queries. A contributing knowledge indicates that every included sub-knowledge is helpful for the predicted SQL, the corresponding content should be valid. Thus, we focus on studying the contribution of generated knowledge to the ground-truth SQL.
@@
where $Y$ is the ground-truth SQL. We check whether every sub-knowledge is contained by $Y$ to judge whether the knowledge contributes. Then, we collect the preference knowledge pairs based on the feedback of ground-truth SQL contribution, formulated as:
@@
This feedback necessitates the generated knowledge to become correctly SQL contributing and reduce the redundancy in the knowledge generation process.
By Eq. 8 and 10, we obtain two preference knowledge set $\mathcal{P}^{db}$ and $\mathcal{P}^{sql}$. To get the final preference learning dataset $\mathcal{D}$, we combine corresponding input of $K_{l}$ according to Eq. 4, denoted by $I_{l}$ and the preference pair $(K_{w},K_{l})$ by: